OpenStudy (anonymous):

Write a proportion for the following problem. The tax on a new car costing $14,400 is $784. At the same tax rate, what is the tax on a vehicle costing $20,000? Round to the nearest dollar. A. 784/14,400 = x/20,000 B. 784/14,400 = 20,000/x C. 20,000/14,400=784/x

OpenStudy (phi):

To do proportions, keep "things together" if they give you the tax and a price, write the ratio tax/price = (different tax)/(different price) or you could write price/tax = (diff price)/(diff tax) BUT NEVER tax/price = (different price)/(different tax) <-- wrong.

OpenStudy (anonymous):

The answer is B?

OpenStudy (phi):

B 784/14,400 = 20,000/x is 784 price or tax is 14,400 price or tax what is 20,000: price or tax?

OpenStudy (anonymous):

784 is price 14,400 is tax and 20,000 is the other price

OpenStudy (anonymous):

yes?

OpenStudy (phi):

yes. Now read this write the ratio tax/price = (different tax)/(different price) or you could write price/tax = (diff price)/(diff tax) BUT NEVER tax/price = (different price)/(different tax) <-- wrong.

OpenStudy (anonymous):

ok so never tax before price

OpenStudy (phi):

oops I lied to you. you said 784 is price 14,400 is tax and 20,000 is the other price but the problem says The tax on a new car costing $14,400 is $784. which means the tax is $784 (generally tax is less than the purchase if you did not know)

OpenStudy (phi):

the car costs 14,400 so that is the price

OpenStudy (phi):

Can you get the answer?

OpenStudy (anonymous):

yes

OpenStudy (phi):

It is A. 784/14,400 = x/20,000